# ZTF SN Ia DR2: Overview

### v1

@@ 5 Data release papers
Data acquisition and processing (light-curve extraction, SALT2 fitting, host matching, redshifts, etc.) are presented in Smith et al., while Lacroix et al. review the accuracy of the DR2 photometry and ongoing work to unlock cosmology-ready calibrations. An overview of the DR2 SN Ia spectra can be found in Johansson et al.
@@
The accuracy of the light-curve modeling is reviewed in Rigault et al. (2024,b ZTFDR2). Additional investigation of light-curve modeling is presented in Kenworthy et al., where we introduce the possibility of an extra stretch-like light curve parameter that may absorb a significant part of the usual phase-independent color term.
@@
Simulation from Amenouche et al. and data distributions presented in Smith et al. and in this overview show that our normal SNe Ia dataset is free from non-random selections at . Based on this volume-limited sample, Ginolin et al. (2024,a ZTFDR2) analyse the SN Ia standardization process and demonstrates the non-linearity of the “bright-slower” relation, while highlighting the most significant environmental magnitude offset to date. The origin of this offset and the color standardization are further discussed in Ginolin et al. (2024,b ZTFDR2) and compared with other higher-redshift samples in Popovic et al. (2024, ZTFDR2). The SN Ia standardization is studied through siblings in Dhawan et al. (2024, ZTFDR2).
@@
Given our large sample statistics, we compare of the SN Ia properties as a function of their cosmic web origin. Ruppin et al. (2024, ZTFDR2) study how SNe Ia vary as a function of their cluster association, notably since cluster galaxies are less star-forming than their field counterparts. Aubert et al. (2024, ZTFDR2) generalize this study when comparing SN Ia properties as a function of their cosmic density field. The actual impact of peculiar motion caused by the velocity field in the derivation of cosmological parameters is discussed in Carreres et al. (2024, ZTFDR2).
@@
Our sample also enables careful studies of SN Ia explosion physics and progenitor origins. Burgaz et al. (2024,a ZTFDR2) presents the spectroscopic diversity (through measurements of key spectral features) of the maximum-light volume-limited sample. This includes investigation of the impact of host-galaxy contamination on spectral measurements, as well as an analysis, including spectral modelling, of the continuum of SN Ia spectral properties seen from the normal SN Ia population to peculiar underluminous subclasses, such as 91bg-like events. Dimitriadis et al. review the photometric diversity of the SN Ia population of the DR2 (including peculiar events), describing how sub-classifications have been derived, as well as intrinsic rates of the sub-classes.
@@
In Terwel et al. we search for late-time interaction with circumstellar material in SNe Ia, finding a few cases of rebrightening years after the explosion. In Harvey et al., we study the demographic of high-velocity silicon features, finding that most SNe Ia exhibit such behavior at early phases and that they are more common in underluminous SNe Ia. Deckers et al. (2024, ZTFDR2) use Gaussian process (GP) fitting of the ZTF light curves to study the properties of the second maximum clearly visible in redder bands, as well as constrain the origin of these features. The connection between host and SNe Ia properties is further studied in Burgaz et al. (b), who investigate the low-mass host SN Ia population, and in Senzel et al., who compare the SN Ia properties on whether they originate from the bulge, bar or disk of the host that is connected to the age and metallicity.
@@
Cosmological analysis will soon follow, once photometric calibrations detailed in Lacroix et al. are completed. This resulting “DR2.5” release will be accompanied by a series of calibration papers. We discourage user to use the current DR2 data to derive cosmological parameters.
@@
| Rigault (b) | Light-curve residuals |
| Kenworthy | Light-curve modeling |
@@ 6 Data Release content and access
The content of this data-release is illustrated in Fig. 7 and the main release parameters are summarize in Table 3. We are providing:
3591 SNe Ia light curves (, , band),
@@
| Light-curve file parameters | |
@@
We do not release light-curve data and derived light-curve properties for the 37 non-peculiar SNe Ia below () from which independent distances could be acquired and that will lead to a measurement of the Hubble constant, . These SNe Ia will be released as part of a dedicated ZTF cosmology program. However, light-curves data of the 12 peculiar SN Ia at SNe Ia are released, as well as spectra and host properties for all targets, including the normal very nearby SNe Ia.
The host tables contains the rest-frame color and stellar mass () both for local (2 kpc radius aperture) and global properties. The host table also contains the host coordinates and SN-host distance information.
The logs table contains pointing (mjd, ra, dec) and observing conditions (zp, limiting magnitude, gain, infobits), sufficient to simulate ZTF DR2 data.
We finally provide additional SN Ia data tables in the tables/extra repository. These are results obtained when changing the used SALT template (SALT2.4, Betoule et al. 2014 or SALT3 Kenworthy et al. 2021, see modeling discussion in Rigault et al. (2024,b ZTFDR2)), or changing the considered phase range (starting at or d, and/or finishing at days) or the use or not of the -band.
All released data can be retrieve from ztfcosmo.in2p3.fr 666data tables available after publication. Our observations are also made available via WISeREP777https://www.wiserep.org(Yaron & Gal-Yam, 2012).
@@ 7 Conclusions
With 3628 objects, we are releasing the largest sample of SN Ia data to date. This release, named “DR2”, publishes all SNe Ia acquired by the ZTF survey up to December 2020 and spectroscopically classified as “Ia”. The classification, as well as the data content have been vetted by multiple members of the science collaboration. Along with 20 companions papers studying in detail SN Ia astrophysics and their use as cosmological distance indicators, we are releasing: 3591 force-photometry light curves (, , ) and SALT2 light curve parameters, 5138 spectra, galaxy and SN-redshifts, global and local (2 kpc) host properties, and observing logs. The SN light curves and light curve properties of normal SNe Ia at are however not released.
This release increases by an order of magnitude current state-of-the-art low-redshift SN Ia dataset. Our sample is homogeneous, with well-controlled selection effects. As part of our release, nearly 1000 SNe Ia are from a volume-limited sample () for which parameter distributions and correlations are representative of the true underlying normal SN Ia population.
The photometric accuracy of the release is at the per-cent level with known per-cent non-linearity. The data are thus of sufficient quality for any internal ZTF SN Ia analysis but are not yet ready to be joined with other datasets to derive cosmological parameters. The Cosmology Science Working group is actively working on final calibration steps, to release the cosmology associated to these SNe Ia. This will constitute the “DR2.5” release that shall become public on an approximately year time-scale. The next step will be the use of the full ZTF survey (2018-2025) that should contain nearly 8,000 spectroscopically confirmed SNe Ia. In addition, we expect 25,000 additional photometrically confirmed SNe Ia. This SNe Ia sample will be our DR3.
@@ Acknowledgements.
- Abdurro’uf et al. (2022) Abdurro’uf, Accetta, K., Aerts, C., et al. 2022, ApJS, 259, 35. doi:10.3847/1538-4365/ac4414
- Astier et al. (2006) Astier, P., Guy, J., Regnault, N., et al. 2006, A&A, 447, 31. doi:10.1051/0004-6361:20054185
- Astier et al. (2013) Astier, P., El Hage, P., Guy, J., et al. 2013, A&A, 557, A55. doi:10.1051/0004-6361/201321668
- Aubert et al. (2024, ZTFDR2) Aubert, M., Rosnet, P., Popovic, B., et al. 2024, arXiv:2406.11680. doi:10.48550/arXiv.2406.11680
- Bellm et al. (2019) Bellm, E. C., Kulkarni, S. R., Graham, M. J., et al. 2019, PASP, 131, 018002. doi:10.1088/1538-3873/aaecbe
- Bellm et al. (2019) Bellm, E. C., Kulkarni, S. R., Barlow, T., et al. 2019, PASP, 131, 068003. doi:10.1088/1538-3873/ab0c2a
- Betoule et al. (2014) Betoule, M., Kessler, R., Guy, J., et al. 2014, A&A, 568, A22. doi:10.1051/0004-6361/201423413
- Blagorodnova et al. (2018) Blagorodnova, N., Neill, J. D., Walters, R., et al. 2018, PASP, 130, 035003. doi:10.1088/1538-3873/aaa53f
- Bloom et al. (2012) Bloom, J. S., Kasen, D., Shen, K. J., et al. 2012, ApJ, 744, L17. doi:10.1088/2041-8205/744/2/L17
- Briday et al. (2022) Briday, M., Rigault, M., Graziani, R., et al. 2022, A&A, 657, A22. doi:10.1051/0004-6361/202141160
- Blondin & Tonry (2007) Blondin, S. & Tonry, J. L. 2007, ApJ, 666, 1024. doi:10.1086/520494
- BOSS Collab. (2017) Alam, S., Ata, M., Bailey, S., et al. 2017, MNRAS, 470, 2617. doi:10.1093/mnras/stx721
- Briday et al. (2022) Briday, M., Rigault, M., Graziani, R., et al. 2022, A&A, 657, A22. doi:10.1051/0004-6361/202141160
- Brout et al. (2019) Brout, D., Sako, M., Scolnic, D., et al. 2019, ApJ, 874, 106. doi:10.3847/1538-4357/ab06c1
- Brout & Scolnic (2021) Brout, D. & Scolnic, D. 2021, ApJ, 909, 26. doi:10.3847/1538-4357/abd69b
- Brout et al. (2022) Brout, D., Scolnic, D., Popovic, B., et al. 2022, ApJ, 938, 110. doi:10.3847/1538-4357/ac8e04
- Burgaz et al. (2024,a ZTFDR2) Burgaz, U., Maguire, K., Dimitriadis, G., et al. 2024, arXiv:2407.06828. doi:10.48550/arXiv.2407.06828
- Carreres et al. (2024, ZTFDR2) Carreres, B., Rosselli, D., Bautista, J. E., et al. 2024, arXiv:2405.20409. doi:10.48550/arXiv.2405.20409
- Childress et al. (2013) Childress, M., Aldering, G., Antilogus, P., et al. 2013, ApJ, 770, 108. doi:10.1088/0004-637X/770/2/108
- Coughlin et al. (2023) Coughlin, M. W., Bloom, J. S., Nir, G., et al. 2023, ApJS, 267, 31. doi:10.3847/1538-4365/acdee1
- Deckers et al. (2022) Deckers, M., Maguire, K., Magee, M. R., et al. 2022, MNRAS, 512, 1317. doi:10.1093/mnras/stac558
- Deckers et al. (2024, ZTFDR2) Deckers, M., Maguire, K., Shingles, L., et al. 2024, arXiv:2406.19460. doi:10.48550/arXiv.2406.19460
- Dekany et al. (2020) Dekany, R., Smith, R. M., Riddle, R., et al. 2020, PASP, 132, 038001. doi:10.1088/1538-3873/ab4ca2
- Desai et al. (2023) Desai, D. D., Kochanek, C. S., Shappee, B. J., et al. 2023, arXiv:2306.11100. doi:10.48550/arXiv.2306.11100
@@
- Dey et al. (2019) Dey, A., Schlegel, D. J., Lang, D., et al. 2019, AJ, 157, 168. doi:10.3847/1538-3881/ab089d
- Dhawan et al. (2022) Dhawan, S., Goobar, A., Smith, M., et al. 2022, MNRAS, 510, 2228. doi:10.1093/mnras/stab3093
- Dhawan et al. (2024, ZTFDR2) Dhawan, S., Mortsell, E., Johansson, J., et al. 2024, arXiv:2406.01434. doi:10.48550/arXiv.2406.01434
- Duev et al. (2019) Duev, D. A., Mahabal, A., Masci, F. J., et al. 2019, MNRAS, 489, 3582. doi:10.1093/mnras/stz2357
- Eisenstein et al. (2005) Eisenstein, D. J., Zehavi, I., Hogg, D. W., et al. 2005, ApJ, 633, 560. doi:10.1086/466512
- Flewelling et al. (2020) Flewelling, H. A., Magnier, E. A., Chambers, K. C., et al. 2020, ApJS, 251, 7. doi:10.3847/1538-4365/abb82d
- Fremling et al. (2020) Fremling, C., Miller, A. A., Sharma, Y., et al. 2020, ApJ, 895, 32. doi:10.3847/1538-4357/ab8943
- Ginolin et al. (2024,a ZTFDR2) Ginolin, M., Rigault, M., Smith, M., et al. 2024, arXiv:2405.20965. doi:10.48550/arXiv.2405.20965
- Ginolin et al. (2024,b ZTFDR2) Ginolin, M., Rigault, M., Copin, Y., et al. 2024, arXiv:2406.02072. doi:10.48550/arXiv.2406.02072
- Goobar & Leibundgut (2011) Goobar, A. & Leibundgut, B. 2011, Annual Review of Nuclear and Particle Science, 61, 251. doi:10.1146/annurev-nucl-102010-130434
- Graham et al. (2019) Graham, M. J., Kulkarni, S. R., Bellm, E. C., et al. 2019, PASP, 131, 078001. doi:10.1088/1538-3873/ab006c
- Gupta et al. (2016) Gupta, R. R., Kuhlmann, S., Kovacs, E., et al. 2016, AJ, 152, 154. doi:10.3847/0004-6256/152/6/154
- Guy et al. (2007) Guy, J., Astier, P., Baumont, S., et al. 2007, A&A, 466, 11. doi:10.1051/0004-6361:20066930
- Guy et al. (2010) Guy, J., Sullivan, M., Conley, A., et al. 2010, A&A, 523, A7. doi:10.1051/0004-6361/201014468
- Hahn et al. (2023) Hahn, C., Wilson, M. J., Ruiz-Macias, O., et al. 2023, AJ, 165, 253. doi:10.3847/1538-3881/accff8
- Holtzman et al. (2008) Holtzman, J. A., Marriner, J., Kessler, R., et al. 2008, AJ, 136, 2306. doi:10.1088/0004-6256/136/6/2306
@@
- Jha et al. (2019) Jha, S. W., Maguire, K., & Sullivan, M. 2019, Nature Astronomy, 3, 706. doi:10.1038/s41550-019-0858-0
- Kasliwal et al. (2019) Kasliwal, M. M., Cannella, C., Bagdasaryan, A., et al. 2019, PASP, 131, 038003. doi:10.1088/1538-3873/aafbc2
- Kelsey et al. (2021) Kelsey, L., Sullivan, M., Smith, M., et al. 2021, MNRAS, 501, 4861. doi:10.1093/mnras/staa3924
- Kenworthy et al. (2021) Kenworthy, W. D., Jones, D. O., Dai, M., et al. 2021, ApJ, 923, 265. doi:10.3847/1538-4357/ac30d8
- Kim et al. (2022) Kim, Y.-L., Rigault, M., Neill, J. D., et al. 2022, PASP, 134, 024505. doi:10.1088/1538-3873/ac50a0
- Le Borgne & Rocca-Volmerange (2002) Le Borgne, D. & Rocca-Volmerange, B. 2002, A&A, 386, 446. doi:10.1051/0004-6361:20020259
- Lezmy et al. (2022) Lezmy, J., Copin, Y., Rigault, M., et al. 2022, arXiv:2209.10882
- Liu et al. (2023) Liu, Z.-W., Röpke, F. K., & Han, Z. 2023, Research in Astronomy and Astrophysics, 23, 082001. doi:10.1088/1674-4527/acd89e
- Maguire et al. (2014) Maguire, K., Sullivan, M., Pan, Y.-C., et al. 2014, MNRAS, 444, 3258. doi:10.1093/mnras/stu1607
- Mannucci et al. (2005) Mannucci, F., Della Valle, M., Panagia, N., et al. 2005, A&A, 433, 807. doi:10.1051/0004-6361:20041411
- Maoz et al. (2014) Maoz, D., Mannucci, F., & Nelemans, G. 2014, ARA&A, 52, 107. doi:10.1146/annurev-astro-082812-141031
- Masci et al. (2019) Masci, F. J., Laher, R. R., Rusholme, B., et al. 2019, PASP, 131, 018003. doi:10.1088/1538-3873/aae8ac
- Müller-Bravo & Galbany (2022) Müller-Bravo, T. & Galbany, L. 2022, The Journal of Open Source Software, 7, 4508. doi:10.21105/joss.04508
- Nicolas et al. (2021) Nicolas, N., Rigault, M., Copin, Y., et al. 2021, A&A, 649, A74. doi:10.1051/0004-6361/202038447
- Nugent et al. (2011) Nugent, P. E., Sullivan, M., Cenko, S. B., et al. 2011, Nature, 480, 344. doi: 10.1038/nature10644
- Papadogiannakis et al. (2019) Papadogiannakis, S., Dhawan, S., Morosin, R., et al. 2019, MNRAS, 485, 2343. doi:10.1093/mnras/stz493
- Perley et al. (2020) Perley, D. A., Fremling, C., Sollerman, J., et al. 2020, ApJ, 904, 35. doi:10.3847/1538-4357/abbd98
- Perlmutter et al. (1999) Perlmutter, S., Aldering, G., Goldhaber, G., et al. 1999, ApJ, 517, 565. doi:10.1086/307221
- Planck Collab. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6. doi:10.1051/0004-6361/201833910
@@
- Popovic et al. (2024, ZTFDR2) Popovic, B., Rigault, M., Smith, M., et al. 2024, arXiv:2406.06215. doi:10.48550/arXiv.2406.06215
- Roman et al. (2018) Roman, M., Hardin, D., Betoule, M., et al. 2018, A&A, 615, A68. doi:10.1051/0004-6361/201731425
- Riess et al. (1996) Riess, A. G., Press, W. H., & Kirshner, R. P. 1996, ApJ, 473, 88. doi:10.1086/178129
- Riess et al. (1998) Riess, A. G., Filippenko, A. V., Challis, P., et al. 1998, AJ, 116, 1009. doi:10.1086/300499
- Riess et al. (2016) Riess, A. G., Macri, L. M., Hoffmann, S. L., et al. 2016, ApJ, 826, 56. doi:10.3847/0004-637X/826/1/56
- Riess et al. (2022) Riess, A. G., Yuan, W., Macri, L. M., et al. 2022, ApJ, 934, L7. doi:10.3847/2041-8213/ac5c5b
- Rigault et al. (2013) Rigault, M., Copin, Y., Aldering, G., et al. 2013, A&A, 560, A66. doi:10.1051/0004-6361/201322104
- Rigault et al. (2015) Rigault, M., Aldering, G., Kowalski, M., et al. 2015, ApJ, 802, 20. doi:10.1088/0004-637X/802/1/20
- Rigault et al. (2019) Rigault, M., Neill, J. D., Blagorodnova, N., et al. 2019, A&A, 627, A115. doi:10.1051/0004-6361/201935344
- Rigault et al. (2020) Rigault, M., Brinnel, V., Aldering, G., et al. 2020, A&A, 644, A176. doi:10.1051/0004-6361/201730404
- Rigault et al. (2024,b ZTFDR2) Rigault, M., Smith, M., Regnault, N., et al. 2024, arXiv:2406.02073. doi:10.48550/arXiv.2406.02073

### v2

@@ 5 Data release papers
Data acquisition and processing (light curve extraction, SALT2 fitting, host matching, redshifts, etc.) are presented in Smith et al. in prep., while Lacroix et al. in prep. review the accuracy of the DR2 photometry and ongoing work to unlock cosmology-ready calibrations. An overview of the DR2 SN Ia spectra can be found in Johansson et al. in prep.
@@
The accuracy of the light curve modeling is reviewed in Rigault et al. (2024,b). An additional investigation of light curve modeling is presented in Kenworthy et al. (2024), in which we introduce the possibility of an extra stretch-like light curve parameter that may absorb a significant part of the usual phase-independent color term.
@@
The simulation from Amenouche et al. (2024) and data distributions presented in Smith et al. in prep. and in this overview show that our normal SNe Ia dataset is free from nonrandom selections at . Based on this volume-limited sample, Ginolin et al. (2024,a) have analysed the SN Ia standardization process and demonstrated the nonlinearity of the “bright-slower” relation, while highlighting the most significant environmental magnitude offset to date. The origin of this offset and the color standardization are further discussed in Ginolin et al. (2024,b) and compared with other higher-redshift samples in Popovic et al. (2024). The SN Ia standardization is studied through siblings in Dhawan et al. (2024).
@@
Given our large sample statistics, we compared the SN Ia properties as a function of their cosmic web origin. Ruppin et al. (2024) study how SNe Ia vary as a function of their cluster association, notably since cluster galaxies are less star-forming than their field counterparts. Aubert et al. (2024) generalize this study when comparing SN Ia properties as a function of their cosmic density field. The actual impact of peculiar motion caused by the velocity field in the derivation of cosmological parameters is discussed in Carreres et al. (2024).
@@
Our sample also enables careful studies of SN Ia explosion physics and progenitor origins. Burgaz et al. (2024,b) present the spectroscopic diversity (through measurements of key spectral features) of the maximum-light volume-limited sample. This includes an investigation of the impact of host-galaxy contamination on spectral measurements, as well as an analysis, including spectral modeling, of the continuum of SN Ia spectral properties seen from the normal SN Ia population to peculiar underluminous subclasses, such as 91bg-like events. Dimitriadis et al. (2024) review the photometric diversity of the SN Ia population of the DR2 (including peculiar events), describing how subclassifications have been derived, as well as intrinsic rates of the subclasses.
@@
In Terwel et al. (2024), we search for late-time interaction with circumstellar material in SNe Ia, finding a few cases of rebrightening years after the explosion. In Harvey et al. (2024), we study the demographic of high-velocity silicon features, finding that most SNe Ia exhibit such behavior at early phases and that they are more common in underluminous SNe Ia. Deckers et al. (2024) use Gaussian process fitting of the ZTF light curves to study the properties of the second maximum clearly visible in redder bands, as well as constrain the origin of these features. The connection between host and SNe Ia properties is further studied in Burgaz et al. (2024,b), who investigate the low-mass host SN Ia population, and in Senzel et al. (2024), who compare the SN Ia properties and whether they originate from the bulge, bar, or disk of the host that is connected to the age and metallicity.
@@
A cosmological analysis will soon follow, once photometric calibrations detailed in Lacroix et al. in prep. are completed. This resulting DR2.5 release will be accompanied by a series of calibration papers. We discourage the user from using the current DR2 data to derive cosmological parameters.
@@
| Rigault (b) | Light curve residuals |
| Kenworthy | Light curve modeling |
@@ 6 Data release content and access
The content of this data release is illustrated in Fig. 7 and the main release parameters are summarized in Table 3. We are providing:
3591 SNe Ia light curves (, , and band),
@@
| Light curve file parameters | |
@@
We have not released light curve data and derived light curve properties for the 37 non-peculiar SNe Ia below () from which independent distances could be acquired and that will lead to a measurement of the Hubble constant, . These SNe Ia will be released as part of a dedicated ZTF cosmology program. However, light curve data of the 12 peculiar SNe Ia at SNe Ia have been released, as well as spectra and host properties for all targets, including the normal very nearby SNe Ia.
The host tables contains the rest-frame color and stellar mass () both for local (2 kpc radius aperture) and global properties. The host table also contains the host coordinates and SN-host distance information. The log table contains pointing (mjd, ra, dec) and observing conditions (zp, limiting magnitude, gain, infobits), which is sufficient to simulate ZTF DR2 data.
We finally provide additional SN Ia data tables in the tables/extra repository. These are results obtained when changing the used SALT template (SALT2.4 Betoule et al. 2014 or SALT3 Kenworthy et al. 2021, see modeling discussion in Rigault et al. (2024,b)), or changing the considered phase range (starting at or d, and/or finishing at days) or the use or not of the band.
All of the released data can be retrieved from ztfcosmo.in2p3.fr.666Data tables available after publication Our observations are also made available via WISeREP (Yaron & Gal-Yam, 2012).777https://www.wiserep.org
@@ 7 Conclusions
With 3628 objects, we are releasing the largest sample of SN Ia data to date. This release, named DR2, publishes all SNe Ia acquired by the ZTF survey up to December 2020 and spectroscopically classified as “Ia.” The classification as well as the data content have been vetted by multiple members of the scientific collaboration. Along with 20 companion papers studying in detail SN Ia astrophysics and their use as cosmological distance indicators, we are releasing: 3591 forced-photometry light curves (, , ) and SALT2 light curve parameters, 5138 spectra, galaxy and SN redshifts, global and local (2 kpc) host properties, and observing logs. The SN light curves and light curve properties of normal SNe Ia at , however, have not been released.
This release increases by an order of magnitude the current state-of-the-art low-redshift SN Ia dataset. Our sample is homogeneous, with well-controlled selection effects. As part of our release, nearly 1000 SNe Ia are from a volume-limited sample () for which parameter distributions and correlations are representative of the true, normal underlying SN Ia population.
The photometric accuracy of the release is at the percent level with known percent nonlinearity. The data are thus of sufficient quality for any internal ZTF SN Ia analysis but are not yet ready to be joined with other datasets to derive cosmological parameters. The Cosmology Science Working group is actively working on final calibration steps, to release the cosmology associated with these SNe Ia. This will constitute the DR2.5 release that will become public on a timescale of approximately one year. The next step will be the use of the full ZTF survey (2018-2025), which should contain nearly 8,000 spectroscopically confirmed SNe Ia. In addition, we expect 25,000 additional photometrically confirmed SNe Ia. This SNe Ia sample will be our DR3.
@@ Acknowledgements.
- Abdurro’uf et al. (2022) Abdurro’uf, Accetta, K., Aerts, C., et al. 2022, ApJS, 259, 35.
- Amenouche et al. (2024) Amenouche, M., Smith, M., Rosnet, P., et al. 2024, arXiv:2409.04650. doi:10.48550/arXiv.2409.04650, (ZTFSI)
- Astier et al. (2006) Astier, P., Guy, J., Regnault, N., et al. 2006, A&A, 447, 31.
- Astier et al. (2013) Astier, P., El Hage, P., Guy, J., et al. 2013, A&A, 557, A55.
- Aubert et al. (2024) Aubert, M., Rosnet, P., Popovic, B., et al. 2024, arXiv:2406.11680. doi:10.48550/arXiv.2406.11680, (ZTFSI)
- Bellm et al. (2019) Bellm, E. C., Kulkarni, S. R., Graham, M. J., et al. 2019, PASP, 131, 018002.
- Bellm et al. (2019) Bellm, E. C., Kulkarni, S. R., Barlow, T., et al. 2019, PASP, 131, 068003.
- Betoule et al. (2014) Betoule, M., Kessler, R., Guy, J., et al. 2014, A&A, 568, A22.
- Blagorodnova et al. (2018) Blagorodnova, N., Neill, J. D., Walters, R., et al. 2018, PASP, 130, 035003.
- Bloom et al. (2012) Bloom, J. S., Kasen, D., Shen, K. J., et al. 2012, ApJ, 744, L17.
- Blondin & Tonry (2007) Blondin, S. & Tonry, J. L. 2007, ApJ, 666, 1024.
- BOSS Collab. (2017) Alam, S., Ata, M., Bailey, S., et al. 2017, MNRAS, 470, 2617.
- Briday et al. (2022) Briday, M., Rigault, M., Graziani, R., et al. 2022, A&A, 657, A22.
- Brout et al. (2019) Brout, D., Sako, M., Scolnic, D., et al. 2019, ApJ, 874, 106.
- Brout & Scolnic (2021) Brout, D. & Scolnic, D. 2021, ApJ, 909, 26.
- Brout et al. (2022) Brout, D., Scolnic, D., Popovic, B., et al. 2022, ApJ, 938, 110.
- Burgaz et al. (2024,a) Burgaz, U., Maguire, K., Dimitriadis, G., et al. 2024, arXiv:2407.06828. (ZTFSI)
- Burgaz et al. (2024,b) Burgaz, U., Maguire, K., Burgaz, U. et al. 2024, A&A, submitteb (ZTFSI)
- Carreres et al. (2024) Carreres, B., Rosselli, D., Bautista, J. E., et al. 2024, arXiv:2405.20409. doi:10.48550/arXiv.2405.20409, (ZTFSI)
- Childress et al. (2013) Childress, M., Aldering, G., Antilogus, P., et al. 2013, ApJ, 770, 108.
- Coughlin et al. (2023) Coughlin, M. W., Bloom, J. S., Nir, G., et al. 2023, ApJS, 267, 31.
- Deckers et al. (2022) Deckers, M., Maguire, K., Magee, M. R., et al. 2022, MNRAS, 512, 1317.
- Deckers et al. (2024) Deckers, M., Maguire, K., Shingles, L., et al. 2024, arXiv:2406.19460. doi:10.48550/arXiv.2406.19460, (ZTFSI)
- Dekany et al. (2020) Dekany, R., Smith, R. M., Riddle, R., et al. 2020, PASP, 132, 038001.
- Desai et al. (2024) Desai, D. D., Kochanek, C. S., Shappee, B. J., et al. 2024, MNRAS, 530, 5016.
@@
- Dey et al. (2019) Dey, A., Schlegel, D. J., Lang, D., et al. 2019, AJ, 157, 168.
- Dhawan et al. (2022) Dhawan, S., Goobar, A., Smith, M., et al. 2022, MNRAS, 510, 2228.
- Dhawan et al. (2024) Dhawan, S., Mortsell, E., Johansson, J., et al. 2024, arXiv:2406.01434. doi:10.48550/arXiv.2406.01434, (ZTFSI)
- Dimitriadis et al. (2024) Dimitriadis, G., Burgaz, U., Deckers, M., et al. 2024, arXiv:2409.04200. doi:10.48550/arXiv.2409.04200, (ZTFSI)
- Duev et al. (2019) Duev, D. A., Mahabal, A., Masci, F. J., et al. 2019, MNRAS, 489, 3582.
- Eisenstein et al. (2005) Eisenstein, D. J., Zehavi, I., Hogg, D. W., et al. 2005, ApJ, 633, 560.
- Flewelling et al. (2020) Flewelling, H. A., Magnier, E. A., Chambers, K. C., et al. 2020, ApJS, 251, 7.
- Fremling et al. (2020) Fremling, C., Miller, A. A., Sharma, Y., et al. 2020, ApJ, 895, 32.
- Ginolin et al. (2024,a) Ginolin, M., Rigault, M., Smith, M., et al. 2024, arXiv:2405.20965. doi:10.48550/arXiv.2405.20965, (ZTFSI)
- Ginolin et al. (2024,b) Ginolin, M., Rigault, M., Copin, Y., et al. 2024, arXiv:2406.02072. doi:10.48550/arXiv.2406.02072, (ZTFSI)
- Goobar & Leibundgut (2011) Goobar, A. & Leibundgut, B. 2011, Annual Review of Nuclear and Particle Science, 61, 251.
- Graham et al. (2019) Graham, M. J., Kulkarni, S. R., Bellm, E. C., et al. 2019, PASP, 131, 078001.
- Gupta et al. (2016) Gupta, R. R., Kuhlmann, S., Kovacs, E., et al. 2016, AJ, 152, 154.
- Guy et al. (2007) Guy, J., Astier, P., Baumont, S., et al. 2007, A&A, 466, 11.
- Guy et al. (2010) Guy, J., Sullivan, M., Conley, A., et al. 2010, A&A, 523, A7.
- Hahn et al. (2023) Hahn, C., Wilson, M. J., Ruiz-Macias, O., et al. 2023, AJ, 165, 253.
- Harvey et al. (2024) Harvey, L., Maguire, K., Burgaz, U. et al. 2024, A&A, submitteb (ZTFSI)
- Holtzman et al. (2008) Holtzman, J. A., Marriner, J., Kessler, R., et al. 2008, AJ, 136, 2306.
@@
- Jha et al. (2019) Jha, S. W., Maguire, K., & Sullivan, M. 2019, Nature Astronomy, 3, 706.
- Kasliwal et al. (2019) Kasliwal, M. M., Cannella, C., Bagdasaryan, A., et al. 2019, PASP, 131, 038003.
- Kelsey et al. (2021) Kelsey, L., Sullivan, M., Smith, M., et al. 2021, MNRAS, 501, 4861.
- Kenworthy et al. (2021) Kenworthy, W. D., Jones, D. O., Dai, M., et al. 2021, ApJ, 923, 265.
- Kenworthy et al. (2024) Kenworthy, W. D., Goobar, A., Jones, D. O. et al. 2024, A&A, submitteb (ZTFSI)
- Kim et al. (2022) Kim, Y.-L., Rigault, M., Neill, J. D., et al. 2022, PASP, 134, 024505.
- Le Borgne & Rocca-Volmerange (2002) Le Borgne, D. & Rocca-Volmerange, B. 2002, A&A, 386, 446.
- Lezmy et al. (2022) Lezmy, J., Copin, Y., Rigault, M., et al. 2022, A&A, 668, A43.
- Liu et al. (2023) Liu, Z.-W., Röpke, F. K., & Han, Z. 2023, Research in Astronomy and Astrophysics, 23, 082001.
- Maguire et al. (2014) Maguire, K., Sullivan, M., Pan, Y.-C., et al. 2014, MNRAS, 444, 3258.
- Mannucci et al. (2005) Mannucci, F., Della Valle, M., Panagia, N., et al. 2005, A&A, 433, 807.
- Maoz et al. (2014) Maoz, D., Mannucci, F., & Nelemans, G. 2014, ARA&A, 52, 107.
- Masci et al. (2019) Masci, F. J., Laher, R. R., Rusholme, B., et al. 2019, PASP, 131, 018003.
- Müller-Bravo & Galbany (2022) Müller-Bravo, T. & Galbany, L. 2022, JOSS, 7, 4508.
- Nicolas et al. (2021) Nicolas, N., Rigault, M., Copin, Y., et al. 2021, A&A, 649, A74.
- Nugent et al. (2011) Nugent, P. E., Sullivan, M., Cenko, S. B., et al. 2011, Nature, 480, 344.
- Papadogiannakis et al. (2019) Papadogiannakis, S., Dhawan, S., Morosin, R., et al. 2019, MNRAS, 485, 2343.
- Perley et al. (2020) Perley, D. A., Fremling, C., Sollerman, J., et al. 2020, ApJ, 904, 35.
- Perlmutter et al. (1999) Perlmutter, S., Aldering, G., Goldhaber, G., et al. 1999, ApJ, 517, 565.
- Planck Collab. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6.
@@
- Popovic et al. (2024) Popovic, B., Rigault, M., Smith, M., et al. 2024, arXiv:2406.06215. doi:10.48550/arXiv.2406.06215, (ZTFSI)
- Roman et al. (2018) Roman, M., Hardin, D., Betoule, M., et al. 2018, A&A, 615, A68
- Riess et al. (1996) Riess, A. G., Press, W. H., & Kirshner, R. P. 1996, ApJ, 473, 88.
- Riess et al. (1998) Riess, A. G., Filippenko, A. V., Challis, P., et al. 1998, AJ, 116, 1009.
- Riess et al. (2016) Riess, A. G., Macri, L. M., Hoffmann, S. L., et al. 2016, ApJ, 826, 56.
- Riess et al. (2022) Riess, A. G., Yuan, W., Macri, L. M., et al. 2022, ApJ, 934, L7.
- Rigault et al. (2013) Rigault, M., Copin, Y., Aldering, G., et al. 2013, A&A, 560, A66.
- Rigault et al. (2015) Rigault, M., Aldering, G., Kowalski, M., et al. 2015, ApJ, 802, 20.
- Rigault et al. (2019) Rigault, M., Neill, J. D., Blagorodnova, N., et al. 2019, A&A, 627, A115.
- Rigault et al. (2020) Rigault, M., Brinnel, V., Aldering, G., et al. 2020, A&A, 644, A176.
- Rigault et al. (2024,b) Rigault, M., Smith, M., Regnault, N., et al. 2024, arXiv:2406.02073. doi:10.48550/arXiv.2406.02073, (ZTFSI)